# CoMPaSS: Enhancing Spatial Understanding in Text-to-Image Diffusion Models

### v1

@@ 5 Conclusion
Table 6: Comparison to state-of-the-art models on the VISOR [gokhale_benchmarking_2023] benchmark. OA stands for “object accuracy”, which is the success rate of the mentioned objects to appear in the image.
Method uncond cond 1 2 3 4 OA
GLIDE [nichol_glide_2022] 1.98 59.06 6.72 1.02 0.17 0.03 3.36
DALLE-mini [Dayma_DALL·E_Mini_2021] 16.17 59.67 38.31 17.50 6.89 1.96 27.10
CogView2 [ding_cogview2_2022] 12.17 65.89 33.47 11.43 3.22 0.57 18.47
Structured Diffusion [feng_training-free_2023] 17.87 62.36 44.70 18.73 6.57 1.46 28.65
DALLE-2 [ramesh_hierarchical_2022] 37.89 59.27 73.59 47.23 23.26 7.49 63.93
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 29.86
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 28.79
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 47.83
SD2.1 +SPRIGHT [chatterjee_getting_2024] 43.23 71.24 71.78 51.88 33.09 16.15 60.68
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 56.95
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 65.56
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 65.78
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 68.23
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 78.64
@@ Appendix A Additional Training Details
Table 7: Hyperparameters used during training.
@@
Training GPUs 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S 2$\times$×L40S
Training Resolution $512\times 512$ $512\times 512$ $512\times 512$ $512\times 512$
@@
A.1 SCOP Dataset Construction
For the SCOP data engine, we empirically determined the optimal threshold values $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ through extensive experimentation. We then applied these thresholds to process the COCO training split [lin_microsoft_2014], resulting in the curated SCOP dataset. The initial Relationship Reasoning stage identifies 2,468,858 object pairs with their corresponding spatial relationships. These pairs are then systematically filtered through the Spatial Constraints Enforcement stage, which applies a series of increasingly stringent criteria: Visual Significance (eliminating 1,929,560 pairs), Semantic Distinction (removing 169,973 pairs), Spatial Clarity (filtering out 119,457 pairs), Minimal Overlap (excluding 148,376 pairs), and Size Balance (removing 73,464 pairs), ultimately yielding 28,028 clear and unambiguous object pairs.
During the training phase, the spatial relationships between object pairs is expressed with one of 8 distinct relationship tokens: <left>, <above>, <right>, <below>, <left+above>, <right+above>, <left+below>, and <right+below>. These relationships are determined based on the exact positions of the objects, with occasional random replacements using the <and> token to enhance robustness. For each pair, we select a square region containing both objects, which is then randomly expanded by up to 10% before being cropped to create the final training image paired with its corresponding text description.
A.2 Model Setups
Our experiments are conducted on four diffusion models, including three UNet-based diffusion models SD1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4, SD1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, SD2.1333https://huggingface.co/stabilityai/stable-diffusion-2-1, and a state-of-the-art MMDiT-based diffusion model FLUX.1-dev [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev.
For the UNet-based models, we discovered that incorporating attention supervision as proposed by [wang_tokencompose_2024] significantly enhanced the convergence of the TENOR module. We therefore integrated this supervision across all UNet-based implementations. In contrast, for the FLUX.1 model, we found that the standard denoising loss alone was sufficient to achieve optimal performance. Notably, our best results were achieved with the CoMPaSS-enhanced FLUX.1 model, using a rank-16 LoRA checkpoint that requires only $\sim$50MiB on-disk storage, making it highly efficient for practical applications.
Detailed training hyperparameters for all model configurations are provided in Tab. 7.
@@ Appendix B Runtime Performance of TENOR
Appendix C More Experimental Results
C.1 Additional Comparisons on VISOR
We present a comprehensive comparison of VISOR metrics against other state-of-the-art models in Tab. 6, demonstrating the superior performance of our approach across various evaluation criteria.
C.2 More Visual Comparisons
To provide a clear visualization of our approach’s effectiveness, we present additional visual comparisons across eight spatial configurations in Figs. 9, 8, 11 and 10. For each prompt, we generate a total of 36 images using different model configurations, offering a comprehensive view of how our method consistently improves spatial understanding across various scenarios and model architectures.
C.3 Ablation Studies on other Models
Table 8: Ablation studies of each component of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Sec. 3.1 of the main paper; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
@@ C.3 Ablation Studies on other Models
While Tab. 4 in the main paper presented ablation studies on SD1.5 and FLUX.1, here we extend our analysis to the other two UNet-based diffusion models SD1.4 and SD2.1 in Tab. 8. The results consistently demonstrate that both components of CoMPaSS contribute to improved spatial understanding across different model architectures.
Figure 7: Example object pairs and their corresponding bounding boxes extracted by the SCOP data engine. Each pair satisfies our spatial constraints for Visual Significance, Semantic Distinction, Spatial Clarity, Minimal Overlap, and Size Balance, ensuring unambiguous spatial relationships.
Figure 8: Additional results demonstrating spatial relationship “left”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 9: Additional results demonstrating spatial relationship “above”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 10: Additional results demonstrating spatial relationship “right”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.
Figure 11: Additional results demonstrating spatial relationship “below”. Each row shows generations from different models for the same text prompt. The first two columns (16 images) are from baseline models (SD1.4, SD1.5, SD2.1, FLUX.1), while the remaining five columns (20 images) are from their CoMPaSS-enhanced counterparts, showing consistent improvements in spatial understanding.

### v2

@@ Appendix A Additional Training Details
A.1 SCOP Dataset Construction
The optimal threshold values for the SCOP data engine were empirically determined to be $\{\tau_{\text{v}},\tau_{\text{u}},\tau_{\text{o}},\tau_{\text{s}}\}=\{0.2,2.0,0.3,0.5\}$ through grid search. We then applied these thresholds to process the COCO training split [43], resulting in the curated SCOP dataset. The initial Relationship Reasoning stage identifies 2,468,858 object pairs with their corresponding spatial relationships. The subsequent Spatial Constraints Enforcement stage systematically filters these pairs. This process applies a series of increasingly stringent criteria: Visual Significance eliminates 1,929,560 pairs, Semantic Distinction removes 169,973, Spatial Clarity filters out 119,457, Minimal Overlap excludes 148,376, and Size Balance removes a final 73,464 pairs. This cascade ultimately yields 28,028 clear and unambiguous object pairs.
During the training phase, the spatial relationships between object pairs are expressed with one of 8 distinct relationship tokens: <left>, <above>, <right>, <below>, <left+above>, <right+above>, <left+below>, and <right+below>. These relationships are determined based on the exact positions of the objects, with occasional random replacements using the <and> token to enhance robustness. For each pair, a square region containing both objects is automatically chosen, which is then randomly perturbed and expanded by up to 10% before being cropped to create the final training image paired with its corresponding text description.
A.2 Model Setups
Table A7: Hyperparameters used during training.
@@ A.2 Model Setups
[…]
Training Resolution $\times 512 512×512$$\times 512$ $\times 512 512×512$$\times 512$
@@
Our experiments are conducted on four diffusion models, including three UNet-based diffusion models SD1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4, SD1.5222https://huggingface.co/stable-diffusion-v1-5/stable-diffusion-v1-5, SD2.1333https://huggingface.co/stabilityai/stable-diffusion-2-1, and the state-of-the-art MMDiT-based diffusion model FLUX.1-dev [^4] https://huggingface.co/black-forest-labs/FLUX.1-dev.
For the UNet-based models, we found that incorporating attention supervision as proposed by [71] significantly enhanced the convergence of the TENOR module. We therefore integrated this supervision across all UNet-based implementations. In contrast, for the FLUX.1 model, the standard denoising loss alone was sufficient to achieve optimal performance. Notably, our best results were achieved with the CoMPaSS-enhanced FLUX.1 model, using a rank-16 LoRA checkpoint that requires only $\sim$50MiB on-disk storage, making it highly efficient for practical applications.
Detailed training hyperparameters for all model configurations are provided in Tab. A7.
@@ Appendix B Runtime Performance of TENOR
Appendix C Additional Results
C.1 Visualization of SCOP Data
We provide visualizations of data curated by SCOP in Fig. A8.
Figure A8: Example object pairs and their corresponding bounding boxes extracted by the SCOP data engine. Each pair satisfies our spatial constraints for Visual Significance, Semantic Distinction, Spatial Clarity, Minimal Overlap, and Size Balance, ensuring unambiguous spatial relationships.
C.2 Additional Comparisons on VISOR
We present a comprehensive comparison of VISOR metrics against other state-of-the-art models in Tab. A10, demonstrating the superior performance of our approach across various evaluation criteria.
C.3 More Visual Comparisons
Figure A9: Additional results demonstrating spatial relationship “left”. Our method consistently improves spatial accuracy over the baseline models.
Figure A10: Additional results demonstrating spatial relationship “above”. Our method consistently improves spatial accuracy over the baseline models.
Figure A11: Additional results demonstrating spatial relationship “right”. Our method consistently improves spatial accuracy over the baseline models.
Figure A12: Additional results demonstrating spatial relationship “below”. Our method consistently improves spatial accuracy over the baseline models.
To provide a clear visualization of our approach’s effectiveness, we present additional visual comparisons across eight spatial configurations in Figs. A10, A9, A12 and A11. For each prompt, we generate a total of 36 images using different model configurations, offering a comprehensive view of how our method consistently improves spatial understanding across various scenarios and model architectures.
C.4 Evaluation on Three-Object Compositions
To further test the generalization capabilities of CoMPaSS, we constructed a more challenging benchmark focused on three-object spatial configurations. This benchmark contains 512 prompts, such as “a clock above a broccoli, with a car to the left of the clock”, and is evaluated using an extension of the GenEval methodology [25]. Despite not being fine-tuned on such complex prompts, CoMPaSS demonstrates a significant improvement in accuracy, as detailed in Tab. A8. This result builds upon the strong two-object performance shown in the main paper (Tab. 2).
Table A8: Multi-object spatial relationship evaluation. Despite being trained only on object pairs, CoMPaSS improves the model’s spatial accuracy on prompts involving three objects.
Model % Correct (any) % Correct (all)
SD2.1 9.1 3.3
SD2.1 +CoMPaSS 32.3 20.8
FLUX.1 30.12 13.1
FLUX.1 +CoMPaSS 52.44 31.2
C.5 Detailed Benchmark Breakdown
While the main paper reported overall scores for standard benchmarks (Tab. 2), here we provide a more granular breakdown. Table A11 presents the full, per-task results for T2I-CompBench, GenEval [25], and DPG-Bench [33], offering a comprehensive performance overview.
C.6 Ablation Studies on other Models
While Tab. 6 in the main paper presented ablation studies on SD1.5 and FLUX.1, here we extend our analysis to the other two UNet-based diffusion models SD1.4 and SD2.1 in Tab. A9. The results consistently demonstrate that both components of CoMPaSS contribute to improved spatial understanding across different model architectures.
Table A9: Ablation study on the components of CoMPaSS. (i) original models; (ii) trained with the SCOP dataset described in Sec. 3.1 of the main paper; (iii) our full method. T2I-CompBench Spatial (T. Spatial) and GenEval Position (G. Pos) scores are reported.
@@ C.6 Ablation Studies on other Models
Table A10: Comparison to state-of-the-art models on the VISOR [26] benchmark. OA stands for “object accuracy”, which measures the rate at which all prompted objects appear in the generated image.
Method uncond cond 1 2 3 4 OA
GLIDE [49] 1.98 59.06 6.72 1.02 0.17 0.03 3.36
DALLE-mini [17] 16.17 59.67 38.31 17.50 6.89 1.96 27.10
CogView2 [19] 12.17 65.89 33.47 11.43 3.22 0.57 18.47
Structured Diffusion [22] 17.87 62.36 44.70 18.73 6.57 1.46 28.65
DALLE-2 [57] 37.89 59.27 73.59 47.23 23.26 7.49 63.93
SD1.4 18.81 62.98 46.60 20.11 6.89 1.63 29.86
SD1.5 17.58 61.08 43.65 18.62 6.49 1.57 28.79
SD2.1 30.25 63.24 64.42 35.74 16.13 4.70 47.83
SD2.1 +SPRIGHT [6] 43.23 71.24 71.78 51.88 33.09 16.15 60.68
FLUX.1 37.96 66.81 64.00 44.18 28.66 14.98 56.95
SD1.4 +CoMPaSS 57.41 87.58 83.23 67.53 49.99 28.91 65.56
SD1.5 +CoMPaSS 61.46 93.43 86.55 72.13 54.64 32.54 65.78
SD2.1 +CoMPaSS 62.06 90.96 85.02 71.29 56.03 35.90 68.23
FLUX.1 +CoMPaSS 75.17 93.22 91.73 83.31 72.21 53.41 78.64
Table A11: Evaluation results of general generation capabilities across a wide range of tasks on GenEval [25], T2I-CompBench [35], and DPG-Bench [33]. While designed to target spatial performance, CoMPaSS also improves overall (Ovr.) alignment scores across most tasks.
Method T2I-CompBench GenEval DPG-Bench
Spat. Col. Shp. Tex. N.Sp. Pos. S.O. T.O. Count Col. Attr. Ovr. Rel. Glb. Ent. Attr. Other Ovr.
SD1.4 0.12 0.38 0.36 0.42 0.31 0.03 0.98 0.41 0.34 0.74 0.06 0.43 81.04 78.12 72.23 72.56 59.60 62.02
SD1.4 +CoMPaSS 0.34 0.49 0.43 0.53 0.31 0.46 0.99 0.68 0.34 0.73 0.17 0.56 83.21 79.33 75.33 72.09 68.00 66.07
SD1.5 0.08 0.38 0.37 0.42 0.31 0.04 0.96 0.38 0.36 0.75 0.06 0.42 73.49 74.63 74.23 75.39 67.81 63.18
SD1.5 +CoMPaSS 0.35 0.50 0.43 0.52 0.31 0.54 0.99 0.69 0.34 0.72 0.15 0.57 84.10 82.67 75.20 73.58 60.80 65.81
SD2.1 0.13 0.51 0.42 0.49 0.31 0.07 0.98 0.51 0.44 0.85 0.17 0.50 83.95 81.16 74.47 75.29 53.60 65.47
SD2.1 +SPRIGHT [6] 0.21 - - - - 0.11 0.99 0.59 0.49 0.85 0.15 0.51 - - - - - -
SD2.1 +CoMPaSS 0.32 0.55 0.43 0.54 0.30 0.51 0.99 0.69 0.20 0.71 0.15 0.54 86.54 79.94 78.89 75.39 62.80 69.48
FLUX.1 0.18 0.69 0.48 0.63 0.31 0.26 0.92 0.77 0.71 0.66 0.27 0.60 92.30 80.55 87.74 85.55 78.40 80.63
FLUX.1 +CoMPaSS 0.30 0.83 0.59 0.71 0.32 0.60 0.99 0.87 0.71 0.80 0.76 0.76 94.12 82.98 90.53 88.30 82.80 84.42